# Mode shapes and sensitivity analysis of torsional vibrations in overhang- and T-shaped microcantilevers

### v1

@@ IV Conclusions
In this study, we analytically derived the frequency characteristic equations and modal sensitivities for the torsional modes of overhang- and T-shaped cantilevers. Our results reveal significant and effective changes in both mode shape and frequency as functions of the overhang length, providing a versatile framework for selecting dimensional parameters to achieve specific performance characteristics. Notably, we presented a detailed analysis of the modal sensitivity of these cantilevers for the first time. Our analytical findings align with previous studies and extend the analysis to overhang-shaped cantilevers. These insights offer valuable guidance for experimentalists in designing cantilever structures with tailored frequencies and sensitivities, enabling highly sensitive measurements across a range of applications.
@@ Acknowledgements
The data in my manuscript can be obtained from the corresponding author upon reasonable request.

### v2

@@ IV Conclusions
In this study, we analytically derived the frequency characteristic equations and modal sensitivities for the torsional modes of overhang- and T-shaped cantilevers. Our results reveal significant and effective changes in both mode shape and frequency as functions of the overhang length, providing a versatile framework for selecting dimensional parameters to achieve specific performance characteristics. Notably, we presented a detailed analysis of the modal sensitivity of these cantilevers for the first time.
By modeling the tip–sample interaction as a linear lateral spring, we introduced a tunable coupling stiffness parameter $\beta_{l}$ that directly affects the torsional response. We showed that increasing $\beta_{l}$ leads to a non-monotonic shift in the modal frequencies, with critical balance points where the torsional frequency of width-varying cantilevers matches that of rectangular ones—offering practical benchmarks for cantilever design. The influence of $\beta_{l}$ on modal sensitivity was found to be strongly geometry-dependent, particularly for overhang-shaped cantilevers where unconventional sensitivity ordering emerged. This behavior suggests that geometric modifications can be strategically used to enhance or suppress sensitivity for specific modes.
Our analytical findings not only align with previous studies for rectangular cantilevers TurnerNanoTech01 , but also extend the analysis to geometries that had not been previously explored in the context of torsional modes. These insights offer valuable guidance for experimentalists in designing cantilever structures with tailored frequencies and sensitivities, enabling highly sensitive measurements across a range of applications. Future work may consider incorporating extended sidewall probes, which have shown promising results in improving dynamic behavior and sensitivity DaiMST07 ; PayamMRT20 .